# Search Arena: Analyzing Search-Augmented LLMs

### v1

@@ Abstract
Search-augmented language models combine web search with Large Language Models (LLMs) to improve response groundedness and freshness. However, analyzing these systems remains challenging: existing datasets are limited in scale and narrow in scope, often constrained to static, single-turn, fact-checking questions. In this work, we introduce Search Arena, a crowd-sourced, large-scale, human-preference dataset of over 24,000 paired multi-turn user interactions with search-augmented LLMs. The dataset spans diverse intents and languages, and contains full system traces with around 12,000 human preference votes. Our analysis reveals that user preferences are influenced by the number of citations, even when the cited content does not directly support the attributed claims, uncovering a gap between perceived and actual credibility. Furthermore, user preferences vary across cited sources, revealing that community-driven platforms are generally preferred and static encyclopedic sources are not always appropriate and reliable. To assess performance across different settings, we conduct cross-arena analyses by testing search-augmented LLMs in a general-purpose chat environment and conventional LLMs in search-intensive settings. We find that web search does not degrade and may even improve performance in non-search settings; however, the quality in search settings is significantly affected if solely relying on the model’s parametric knowledge. We open-sourced the dataset to support future research in this direction.
[…]
@@ 1 Introduction
Large Language Models (LLMs) have become a popular interface for human–AI interaction, supporting information seeking and task assistance through natural, multi-turn dialogue. However, the capabilities of these models are constrained by their reliance on static training data, which prevents them from effectively handling time-sensitive questions, emerging topics, or niche domains. Search-augmented LLMs aim to bridge this gap by retrieving and using live web data during inference. Access to search enables LLMs to provide up-to-date, domain-specific, and factually verifiable responses [24, 33]. Recent developments [16, 41, 47] also reflect the growing interest in search-augmented LLMs.
Despite rapid progress in developing search-augmented LLMs, our understanding of how users interact with these systems—what they ask, how they engage in multi-turn dialogue, and what they expect in return—remains limited. Existing datasets capture interactions with either standalone LLMs [10, 70, 71] or traditional web search engines [8, 13]. However, search-augmented LLMs represent a hybrid interface different from both: they not only retrieve information through web search but also rely on their reasoning and conversational capabilities. The most widely used datasets for evaluating these systems (SimpleQA [60] and BrowseComp [61]) primarily consist of single-turn, monolingual, fact-based queries and are relatively small in scale (typically $\leq$5k queries). As shown in Figure 1, fact-checking accounts for only one-fifth of real-world user queries; the majority of user prompts, such as seeking analyses, recommendations, or problem-solving guidance, require a combination of factual retrieval, reasoning, and open-ended dialogue. User expectations also extend beyond factual correctness: preferences can be shaped by the number, relevance, and credibility of citations, as well as the presentation style of responses.
Table 1: Comparison of Search Datasets. Unlike prior datasets such as SimpleQA [60] and BrowseComp [61], which consist of static, English-only, single-turn fact-seeking queries, Search Arena evaluates models in diverse, open-ended, multilingual, and multi-turn settings. We release 24,069 conversations including 12,652 human preference votes. Further analyses are provided in Section 2.
@@
Figure 1: (Left) Nine intent categories with representative examples (truncated). In-the-wild user prompts are often ambiguous and require real-time web retrieval. (Right) Distribution of intents across user prompts. The majority of queries require more than a simple factual lookup and range from information synthesis to creative content generation. The Other category is excluded from the visualizations.
@@
We not only analyze user prompts to search-augmented LLMs, but also their preferences. We model user preferences through the Bradley–Terry model [5, 10, 55] and study how different response characteristics interact with user judgments. We find that reasoning, larger search context window, and longer responses are positively associated with user preferences. Since citations are central to the trustworthiness of web-grounded outputs, we also examine citation features. Our results show that users prefer responses with a higher number of cited sources (Figure 4). In addition, users prefer model responses citing tech-related platforms, community blogs, and social networks, but less favor Wikipedia (Figure 6). While correctly attributed citations, as expected, positively interact with user preferences ($\beta=0.285$), we also observe a positive association between user preference and the number of irrelevant citations ($\beta=0.273$). This finding raises concerns that users may be overly influenced by the presence of citations, even when they do not support the associated claims.
[…]
@@ 2 Human Preference Dataset in Search
We launched Search Arena, an open, crowd-sourced evaluation platform for search-augmented LLMs, on March 18, 2025. Search Arena is implemented as a separate tab within the Chatbot Arena web application [10], where users interact exclusively with search-enabled models. The search mode interface encourages more search-intensive queries, as users adjust their expectations compared to the default Text Arena interface. During each session, two anonymous models respond to a user query, the user can then cast a vote for their preferred model response. Details on user interface and potential limitations are reported in Appendix A and Section 5.
Between March 18 and May 8, we collected more than 24,000 conversations and 12,000 user votes across 13 models, spanning a range of model configurations (e.g., reasoning models, search context sizes, etc.). We release the full dataset, with each record containing model identities, the user vote, conversation histories, and system metadata (e.g., reasoning traces, retrieved URLs). Table 1 presents key differences between Search Arena and prior datasets (SimpleQA [60] and BrowseComp [61]), including dataset scale, prompt characteristics, and available metadata. In the following subsections, we analyze prompt distributions across linguistic and intent dimensions in comparison to existing benchmark datasets.
Figure 2: (Left) Search Arena prompt language distribution. The dataset is multilingual, spanning over 70 languages, with English prompts accounting for 58.3% of the data. (Right) Prompt length distribution of Search Arena (blue), BrowseComp (purple), and SimpleQA (green). Search Arena prompt lengths are more spread out and cover the range of BrowseComp [61] and SimpleQA [60] questions.
@@ 2.2 Intent Diversity
Existing search-augmented LLM evaluation datasets focus solely on factuality. To study how in-the-wild user prompts from Search Arena differ from SimpleQA [60] and BrowseComp [61] questions, we apply an LLM-based dataset differencing framework [14, 72]. We also compare Search Arena prompts with the prompt distribution of Text Arena [10], where users’ expectations of the models are not influenced by search settings. GPT-4.1 is used for generation and GPT-4.1-mini for ranking of candidate distinguishing properties; more details on the pipeline are provided in Appendix C. The summaries of the top properties reveal high-level pairwise differences between the datasets:
@@
These summaries show high-level differences between prompt distributions across three settings–user prompts to search-augmented LLMs (Search Arena), static factuality questions (SimpleQA and BrowseComp), and prompts to regular LLMs (Text Arena). We also study Search Arena prompts through topic modeling and observe diverse and real-time topics, ranging from market analysis to health discussions (see Figure A7).
For a more structured and in-depth analysis, we introduce a taxonomy of user intent categories. While prior work has explored user intent classification in general dialogue settings [37, 50], our work focuses on real-world interactions in search-augmented chat-based settings. The taxonomy includes nine categories: Factual Lookup, Information Synthesis, Analysis, Recommendation, Explanation, Creative Generation, Guidance, Text Processing, and Other. We use secondary labels for ambiguous or multi-purpose prompts. Details on the taxonomy design process, as well as descriptions of categories, are provided in Appendix B. We scale the annotation to the full dataset using GPT-4.1, with a manually tuned prompt seeded on 100 examples and validated on 150 multilingual prompts. The resulting Cohen’s kappa of 0.812 indicates strong agreement between model- and human-annotated labels. The annotation pipeline and validation details are provided in Appendix B.
@@ 3 Preference Analyses in Search
With over 12,000 anonymized human preference votes, Search Arena supports fine-grained analysis of how different response features interact with user preferences. We report model performance through both head-to-head win rates as well as scaled coefficients estimated using the Bradley–Terry model [5, 10] in Appendix D. To analyze how features interact with user preferences, we follow prior work [55] by adding normalized differences between pairwise features to the Bradley-Terry model and reporting the fitted coefficients.
In this section, we first focus on two key feature groups: (1) general features such as model type, search depth, and response length (Subsection 3.1), and (2) citation-related features, such as number of citations, citation sources, and citation attributions (Subsection 3.2). We then study how search and non-search models generalize in search-heavy and general chat scenarios in Subsection 3.3. Additional details and further analyses are provided in Appendix D and Appendix E.
@@ 3.1 General Features
Reasoning. Reasoning models generally perform better under the Search Arena prompt distribution, with the top three reasoning models achieving over 60.0% average win rates, suggesting that reasoning improves performance in search-augmented chat interactions. Consistent with prior work [15], we observe prompt analysis, problem decomposition, and backtracking behaviors in reasoning traces. Additionally, we find that reasoning models not only interpret and analyze retrieved content but also rerank sources and filter out irrelevant information. One example of a reasoning trace is in Figure 3 (Left).
@@
Response Length. Consistent with findings from prior work [10, 52, 55], we observe that users are biased towards more verbose responses. The Bradley-Terry coefficient corresponding to response length is positive and statistically significant ($\beta_{\text{length}}=0.334$), indicating that users tend to prefer longer answers. The positive correlation between model score and response length is also shown in Figure 4 (Left). Additionally, Figure 5 (Left) shows the length distribution across eight user intent categories; responses to Factual Lookup prompts are much shorter (168.3 words on average) compared to Creative Generation (422.8 words) and Analysis (393.2 words) prompts. Furthermore, we find that the response length coefficient on Factual Lookup prompts ($\beta_{\text{length, factual}}=0.156$) is 2.14 times smaller than the effect on the full dataset, suggesting that users prefer less verbose answers to Factual Lookup queries compared to other categories.
@@ 3.2 Citation Features
Citations are central to the trustworthiness of web-grounded outputs in search-enabled scenarios and are a unique feature in Search Arena compared to the prior Text Arena [10, 70]. Therefore, we further examine how these factors influence user preference along three dimensions: number of cited sources, types of cited sources, and attribution of inline citations to the generated content.
@@
Types of Cited Sources. We categorize domains of retrieved URLs into nine groups (e.g., news, Wikipedia, social media, tech/code platforms, etc.); categorization details are provided in Appendix E. Figure 6 (Left) shows source category coefficient estimates with 95% confidence intervals; we also control for the number of citations to account for the bias shown in the previous section. Citing tech-related platforms (e.g., Stack Overflow), community platforms (e.g., Substack), and social media (e.g., TikTok) are positively associated with user preferences with fitted coefficients equal to $\beta_{\text{tech}}=0.073$, $\beta_{\text{community}}=0.061$, and $\beta_{\text{social}}=0.057$, respectively. Surprisingly, citing Wikipedia is negatively correlated with user preferences ($\beta_{\text{wiki}}=-0.071$). To interpret the latter result, we inspect rejected model responses citing Wikipedia and identify two potential explanations: (1) Wikipedia articles are often very lengthy and broad, not directly relevant to a user’s question, and (2) citing Wikipedia is not preferred for queries requiring real-time information. A qualitative example is shown in Figure 3 (Right).
@@
Citation Attribution. We then study how the correctness of citation-to-claim attribution (i.e., whether the inline citation supports the attributed claim) interacts with user preferences. Formally, for each multi-turn interaction, we decompose model responses into a set of claim-citation pairs ${(c_{i},\;u_{i})}$, where $c_{i}$ denotes a textual claim, and $u_{i}$ is the corresponding inline citation. For each pair, we evaluate whether the webpage content $D_{i}$ supports, is irrelevant to, or contradicts the claim $c_{i}$. This process is automated via an LLM-based pipeline described in Figure 7 through an example. Due to scraping challenges and the high cost of LLM calls, we run the pipeline on roughly 100 English conversations per intent category. The resulting output of each conversation is a set of triplet $\{(c_{i},u_{i},t_{i})\}_{i=1}^{N}$, where $t_{i}\in\{\text{Support},\text{Irrelevant},\text{Contradict}\}$ and $N$ is the total number of claims per conversation. We then compute the number of supporting, irrelevant, and contradicting claims per model response and add them as control covariates in the Bradley–Terry analysis. Implementation details, including scraping tools, parsing logic, and validation process, are provided in Appendix E.

### v2

# arXiv:2506.05334v2[cs.CL] 02 Mar 2026

@@ Abstract
Search-augmented language models combine web search with Large Language Models (LLMs) to improve response groundedness and freshness. However, analyzing these systems remains challenging: existing datasets are limited in scale and narrow in scope, often constrained to static, single-turn, fact-checking questions. In this work, we introduce Search Arena, a crowd-sourced, large-scale, human-preference dataset of over 24,000 paired multi-turn user interactions with search-augmented LLMs. The dataset spans diverse intents and languages, and contains full system traces with around 12,000 human preference votes. Our analysis reveals that user preferences are influenced by the number of citations, even when the cited content does not directly support the attributed claims, uncovering a gap between perceived and actual credibility. Furthermore, user preferences vary across cited sources, revealing that community-driven platforms are generally preferred and static encyclopedic sources are not always appropriate and reliable. To assess performance across different settings, we conduct cross-arena analyses by testing search-augmented LLMs in a general-purpose chat environment and conventional LLMs in search-intensive settings. We find that web search does not degrade and may even improve performance in non-search settings; however, the quality in search settings is significantly affected if solely relying on the model’s parametric knowledge. We open-sourced the dataset to support future research.
[Uncaptioned image] Search Arena Code [Uncaptioned image] Dataset
@@ 1 Introduction
Large Language Models (LLMs) have become a popular interface for human–AI interaction, supporting information seeking and task assistance through natural, multi-turn dialogue. However, the capabilities of these models are constrained by their reliance on static training data, which prevents them from effectively handling time-sensitive questions, emerging topics, or niche domains. Search-augmented LLMs aim to bridge this gap by retrieving and using live web data during inference. Access to search enables LLMs to provide up-to-date, domain-specific, and factually verifiable responses (Hilton et al., 2021; Li et al., 2023a). Recent developments (Gemini, 2024; OpenAI, 2024; Perplexity, ) also reflect the growing interest in search-augmented LLMs.
Despite rapid progress in developing search-augmented LLMs, our understanding of how users interact with these systems—what they ask, how they engage in multi-turn dialogue, and what they expect in return—remains limited. Existing datasets capture interactions with either standalone LLMs (Chiang et al., 2024; Zhao et al., 2024; Zheng et al., 2024) or traditional web search engines (Chen et al., 2024b; Craswell et al., 2020). However, search-augmented LLMs represent a hybrid interface different from both: they not only retrieve information through web search but also rely on their reasoning and conversational capabilities. The most widely used datasets for evaluating these systems (SimpleQA (Wei et al., 2024) and BrowseComp (Wei et al., 2025)) primarily consist of single-turn, monolingual, fact-based queries and are relatively small in scale (typically $\leq$5k queries). As shown in Figure 1, fact-checking accounts for only one-fifth of real-world user queries; the majority of user prompts, such as seeking analyses, recommendations, or problem-solving guidance, require a combination of factual retrieval, reasoning, and open-ended dialogue. User expectations also extend beyond factual correctness: preferences can be shaped by the number, relevance, and credibility of citations, as well as the presentation style of responses.
Table 1: Comparison of Search Datasets. Unlike prior datasets such as SimpleQA Wei et al. (2024) and BrowseComp Wei et al. (2025), which are static, English-only, single-turn fact-seeking queries, Search Arena evaluates models in diverse, open-ended, multilingual, and multi-turn settings. We release 24,069 conversations with 12,652 preference votes. Further analyses are provided in Section 2.
@@
Figure 1: (Left) Nine intent categories with representative examples (truncated). In-the-wild prompts are often ambiguous and require real-time web retrieval. (Right) Intent distribution across prompts. Most queries require more than a simple fact lookup, ranging from information synthesis to creative generation. The Other category is excluded.
@@
We not only analyze user prompts to search-augmented LLMs, but also their preferences. We model user preferences through the Bradley-Terry model (Bradley & Terry, 1952; Chiang et al., 2024; Tianle Li, 2024) and study how different response characteristics interact with user judgments. We find that reasoning, a larger search context window, and longer responses are positively associated with user preferences. Since citations are central to trustworthy web-grounded outputs, we also examine citation features. Our results show that users prefer responses with more cited sources (Figure 4) and those citing tech-related platforms, community blogs, and social networks, but less Wikipedia (Figure 6). While correctly attributed citations positively interact with preferences ($\beta=0.285$), we observe a positive association with irrelevant citations ($\beta=0.273$). This raises concerns that users may be overly influenced by citation presence, even when they do not support the associated claims.
We also investigate how search-augmented and non-search models perform across different settings by deploying a non-search LLM in the Search Arena, and a search-augmented LLM in a general-purpose Text Arena. Our results show that conventional LLMs underperform in search-intensive settings ($\text{p-value}=0.009$). However, search-augmented models perform comparably in general chat settings, with improved performance on factual lookup queries ($\text{p-value}=0.012$) and slightly degraded performance on text processing prompts ($\text{p-value}=0.077$).
@@ 2 Human Preference Dataset in Search
We launched Search Arena, an open, crowd-sourced evaluation platform for search-augmented LLMs, on March 18, 2025. The Search Arena is implemented as a separate tab within the Chatbot Arena web application (Chiang et al., 2024), where users interact exclusively with search-enabled models. The search mode interface encourages more search-intensive queries, as users adjust their expectations. During each session, two anonymous models respond to a user query, the user can then cast a vote for their preferred model response. Details on user interface and potential limitations are reported in Appendix A and Section 5.
Between March 18 and May 8, we collected more than 24,000 conversations and 12,000 user votes across 13 models, spanning a range of model configurations (e.g., reasoning models, search context sizes, etc.). The collected dataset contains model identities, the user vote, conversation histories, and system metadata (e.g., reasoning traces, retrieved URLs). Table 1 presents key differences between Search Arena and prior datasets (SimpleQA (Wei et al., 2024) and BrowseComp (Wei et al., 2025)), including dataset scale, prompt characteristics, and available metadata. In the following subsections, we analyze prompt distributions across linguistic and intent dimensions in comparison to existing benchmark datasets.
Figure 2: (Left) Search Arena prompt language distribution. The dataset is multilingual, spanning over 70 languages, with English prompts accounting for 58.3% of the data. (Right) Prompt length distribution of Search Arena (blue), BrowseComp (purple), and SimpleQA (green). Search Arena prompt lengths are more spread out and cover the range of BrowseComp Wei et al. (2025) and SimpleQA Wei et al. (2024) questions.
@@ 2.2 Intent Diversity
Existing search-augmented LLM evaluation datasets focus solely on factuality. To study how in-the-wild user prompts from Search Arena differ from SimpleQA (Wei et al., 2024) and BrowseComp (Wei et al., 2025) questions, we apply an LLM-based dataset differencing framework (Dunlap et al., 2024; Zhong et al., 2022). We also compare Search Arena prompts with the prompt distribution of Text Arena (Chiang et al., 2024), where users’ expectations of the models are not influenced by search settings. GPT-4.1 is used for generation and GPT-4.1-mini for ranking of candidate distinguishing properties; more details on the pipeline are provided in Appendix C. The summaries of the top properties reveal high-level pairwise differences between the datasets:
@@
These summaries show high-level differences between prompt distributions across three settings–user prompts to search-augmented LLMs (Search Arena), static factuality questions (SimpleQA and BrowseComp), and prompts to regular LLMs (Text Arena). We also study Search Arena prompts through topic modeling and observe diverse and real-time topics, ranging from market analysis to health discussions (see Figure B.4).
For a structured and in-depth analysis, we introduce a taxonomy of user intent categories. While prior work has explored user intent classification in general dialogue settings (Liu et al., 2024b; Shah et al., 2023), we focus on real-world interactions in search-augmented chat-based settings. The taxonomy includes nine categories: Factual Lookup, Information Synthesis, Analysis, Recommendation, Explanation, Creative Generation, Guidance, Text Processing, and Other. We use secondary labels for ambiguous or multi-purpose prompts. Details on the taxonomy design, as well as descriptions of categories, are provided in Appendix B. We scale the annotation to the full dataset using GPT-4.1, with a manually tuned prompt seeded on 100 examples and validated on 150 multilingual prompts. The resulting Cohen’s kappa of 0.812 indicates strong agreement between model- and human-annotated labels. The annotation pipeline and validation details are provided in Appendix B.
@@
We report comparative analysis with additional related datasets (CORAL, WildChat) in Appendix G.
@@ 3 Preference Analyses in Search
With over 12,000 anonymized human preference votes, Search Arena supports fine-grained analysis of how different response features interact with user preferences. We report model performance through both head-to-head win rates as well as scaled coefficients estimated using the Bradley-Terry model (Bradley & Terry, 1952; Chiang et al., 2024) in Appendix D. To analyze how features interact with user preferences, we follow prior work (Tianle Li, 2024) by adding normalized differences between pairwise features to the Bradley-Terry model and reporting the fitted coefficients.
As the correlational analysis in this section relies on user-provided votes, we first clarify their subjective nature and assess their reliability. We do not treat the human-preferred response as the objectively “correct” one, but rather as the response better aligned with the user’s personal preferences. In our evaluation and data collection setup, the prompt is tied to the judge (i.e., user): this pairing is critical, as only the original user knows the hidden intent behind their prompt. To verify that users are not voting randomly, we conducted a small study on 100 samples with 3 expert annotators and found expert–user agreement to be 68% when excluding tie votes (random agreement = 50%). Given the inherent subjectivity and the lack of a single “correct” answer for many prompts, some disagreement is expected.
In the following section, our analysis is purely correlational rather than causal, and we view more controlled causal investigation as an important direction for future work. We first focus on two key feature groups: (1) general features such as model type, search depth, and response length (Subsection 3.1), and (2) citation-related features, such as number of citations, citation sources, and citation attributions (Subsection 3.2). We then study how search and non-search models generalize in search-heavy and general chat scenarios in Subsection 3.3. Additional details and further analyses are provided in Appendix D and Appendix E.
@@ 3.1 General Features
Reasoning. Reasoning models generally perform better in Search Arena prompt distribution, with the top three models achieving over 60.0% average win rates, suggesting that reasoning improves performance in search-augmented chat interactions. Consistent with prior work (Gandhi et al., 2025), we observe prompt analysis, problem decomposition, and backtracking behaviors in reasoning traces. Additionally, we find that reasoning models not only interpret and analyze retrieved content but also rerank sources and filter out irrelevant information. One example is in shown Figure 3 (Left).
@@
Response Length. Consistent with findings from prior work (Chiang et al., 2024; Steyvers et al., 2024; Tianle Li, 2024), we observe that users are biased towards more verbose responses. The Bradley-Terry coefficient corresponding to response length is positive and statistically significant ($\beta_{\text{length}}=0.334$), indicating that users tend to prefer longer answers. The positive correlation between model score and response length is also shown in Figure 4 (Left). Additionally, Figure 5 (Left) shows the length distribution across eight user intent categories; responses to Factual Lookup prompts are much shorter (168.3 words on average) compared to Creative Generation (422.8 words) and Analysis (393.2 words) prompts. Furthermore, we find that the response length coefficient on Factual Lookup prompts ($\beta_{\text{length, factual}}=0.156$) is 2.14 times smaller than the effect on the full dataset, suggesting that users prefer less verbose answers to Factual Lookup queries compared to other categories.
@@ 3.2 Citation Features
Citations are central to the trustworthiness of web-grounded outputs in search-enabled scenarios and a unique feature of Search Arena compared to Text Arena (Chiang et al., 2024). Therefore, we further examine how these factors influence user preference along three dimensions: number of cited sources, types of cited sources, and attribution of inline citations.
@@
Types of Cited Sources. We categorize retrieved URL domains into nine groups (e.g., news, Wikipedia, social media, tech/code platforms, etc.); categorization details appear in Appendix E. Figure 6 (Left) shows source category coefficient estimates with 95% confidence intervals; we also control for citation counts to account for the bias shown earlier. Citing tech-related platforms (e.g., Stack Overflow), community platforms (e.g., Substack), and social media (e.g., TikTok) are positively associated with user preferences with fitted coefficients equal to $\beta_{\text{tech}}=0.073$, $\beta_{\text{community}}=0.061$, and $\beta_{\text{social}}=0.057$, respectively. Surprisingly, citing Wikipedia is negatively correlated with user preferences ($\beta_{\text{wiki}}=-0.071$). To interpret the latter result, we inspect rejected model responses citing Wikipedia and identify two potential explanations: (1) Wikipedia articles are often very lengthy and broad, not directly relevant to a user’s question, and (2) citing Wikipedia is not preferred for queries requiring real-time information. A qualitative example appears in Figure 3 (Right).
@@
Citation Attribution. We then study how the correctness of citation-to-claim attribution (i.e., whether the inline citation supports the attributed claim) interacts with user preferences. Formally, for each multi-turn interaction, we decompose model responses into a set of claim-citation pairs $(c_{i},\;u_{i})$}, where $c_{i}$ denotes a textual claim, and $u_{i}$ is the corresponding inline citation. For each pair, we evaluate whether the webpage content $D_{i}$ supports, is irrelevant to, or contradicts the claim $c_{i}$. This process is automated via an LLM-based pipeline described in Figure 7 through an example. Due to scraping challenges and the high cost of LLM calls, we run the pipeline on roughly 100 conversations per intent category (800 examples in total). The resulting output of each conversation is a set of triplet $\{(c_{i},u_{i},t_{i})\}_{i=1}^{N}$, where $t_{i}\in\{\text{Support},\text{Irrelevant},\text{Contradict}\}$ and $N$ is the total number of claims per conversation. We then compute the number of supporting, irrelevant, and contradicting claims per model response and add them as control covariates in the Bradley-Terry analysis. Implementation details, including scraping tools, parsing logic, and validation process, are provided in Appendix E.